Solve the following linear equations: z + 8 = 9 Falling Behind in Studies? Join Our Performance Improvement Batch

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Dear Student z + 8 = 9 z = 9 -8 z =1 Thanks